# MedCLIP-SAMv2: Towards Universal Text-Driven Medical Image Segmentation

### v3

@@ 4.3.1 Effect of text prompt designs
We conducted a series of experiments to analyze the impact of various text prompt designs on zero-shot segmentation performance. In particular, we compared six different prompt configurations: P0 and P1 include the class name of the object to be segmented, while P2 and P3 consist of longer, descriptive single prompts, and finally P4 and P5 are ensembles of 20 text prompts. Note that P0, P2, and P4 are generic text prompts, while P1, P3, and P5 are more nuanced with subtypes of the target object of interests. For example, for Breast Ultrasound, P0 is “breast tumor” while P1 can either be “malignant breast tumor” or “benign breast tumor” depending on the tumor class. For P2, we used one descriptive sentence, such as “A medical breast mammogram showing a suspicious, irregularly shaped mass suggestive of a breast tumor.P3, on the other hand, includes descriptive text about a specific tumor subtype, like “A medical breast mammogram showing an irregularly shaped, spiculated mass suggestive of a malignant breast tumor.P4 and P5 are similar to P2 and P3, but they use an ensemble approach by averaging the text embeddings of 20 different prompts. Here, all descriptive clinical prompts are generated using GPT-4 [55]. For Lung CT, we evaluated solely on generic prompts as there is only one class available. As shown in Table 3, the choice of text prompt significantly influences segmentation performance. Class-specific prompts (P3) generally yielded better results for smaller structures like breast and brain tumors whereas generic prompts (P0, P2) performed better for larger structures like lungs in X-ray and CT scans, where simpler, more generic descriptions allowed the model to focus on larger areas. The best prompt configuration for each task is used to generate the results presented in Table 1.
Prompt Breast Ultrasound Brain MRI Lung X-ray Lung CT
P0 65.0415.84 69.3816.14 72.008.47 77.788.48 76.392.14 81.632.15 71.146.02 73.175.95
P1 68.8815.03 73.0215.39 38.6811.83 41.3712.33 71.035.31 75.075.33 - -
P2 70.2612.98 74.6013.35 72.517.92 78.517.85 65.235.54 69.085.68 80.925.51 82.765.44
P3 78.429.32 82.699.50 77.126.82 83.186.70 65.355.45 69.295.55 - -
P4 68.8617.18 72.3117.67 70.469.04 75.699.25 70.225.46 74.075.53 77.385.76 79.195.66
P5 66.4118.18 69.9918.71 71.068.68 76.468.73 69.825.31 73.785.39 - -
Table 3: Effect of different text prompt templates on the segmentation performance (meanstd)
4.3.2 Component Ablation Analysis
@@
Table 4 shows the contribution of each component of our framework in improving the average segmentation performance on all datasets. Starting with saliency maps generated using the M2IB, we achieved a baseline DSC of 46.23 and an NSD of 50.50, providing an initial focus on key regions of interest. Fine-tuning BiomedCLIP with the proposed DHN-NCE loss raised the DSC to 49.43 and the NSD to 53.54. Post-processing the saliency maps with k-means clustering further enhanced the segmentation quality, allowing the model to better delineate foreground and background areas by refining the initial segmentation boundaries. Incorporating a filtering step greatly impacted the results, increasing the DSC to 58.80 and the NSD to 63.10, as it eliminated small, irrelevant clusters and reduced noise, improving overall precision. With the integration of SAM and the use of visual prompts, such as bounding boxes or points, our zero-shot method yielded a substantial improvement, achieving a DSC of 78.21 and an NSD of 82.57. Finally, weakly supervised training with checkpoint ensembling further refined the segmentation quality by leveraging pseudo-labels generated from the zero-shot method. By using these pseudo-labels to fine-tune a segmentation network, we were able to reach a final DSC of 82.11 and an NSD of 87.33.
@@
1: Saliency Maps 46.236.58 50.506.86
2: + DHN-NCE Fine-tuning 49.105.46 53.545.62
3: + K-Means Post-processing 52.144.27 55.894.40
4: + Area-based Filtering 58.804.73 63.104.95
5: + SAM 78.215.70 82.575.55
6: + nnUNet Ensemble 82.117.77 87.337.70
Table 4: Effect of different components (meanstd)
4.3.3 Effect of Saliency Maps
As shown in Table 5, M2IB achieved the highest performance across all tasks, with an average DSC of 78.21 and NSD of 82.57 when using the fine-tuned BiomedCLIP model. In both its pre-trained and fine-tuned forms, M2IB significantly outperformed gScoreCAM and GradCAM ($p$ $<$ 0.05). BiomedCLIP fine-tuning improved the scores across all saliency map techniques on average, with the largest gains seen in M2IB, which improved by 3.98 in DSC and 4.10 in NSD compared to its pre-trained version.
@@ 4.3.3 Effect of Saliency Maps
Pre-trained BiomedCLIP M2IB 74.236.14 78.476.02
gScoreCAM 59.866.02 62.965.86
GradCAM 28.788.59 31.878.34
Fine-tuned BiomedCLIP M2IB 78.215.70 82.575.55
gScoreCAM 61.736.22 64.996.09
GradCAM 30.608.90 33.818.73
Table 5: Comparison between different Saliency Map techniques as well as the pre-trained and fine-tuned BiomedCLIP on the overall performance (meanstd)
Figure 2: Qualitative comparison of segmentation results. Coarse=K-means post-processed saliency map, WSS=Weakly Supervised Segmentation and GT=Ground Truth. The uncertainty map corresponds to the weakly supervised segmentation.
@@
SAM ViT-H Points 66.305.76 69.855.90 66.904.69 71.924.83 76.392.14 81.632.15 62.115.25 64.635.28
BBoxes 78.429.32 82.699.50 77.126.82 83.186.70 72.404.72 76.384.76 80.925.51 82.765.44
Points + BBoxes 75.645.57 79.605.62 76.043.66 81.763.64 76.221.94 81.301.93 63.844.72 66.234.71
SAM-Med2D ViT-B Points 74.525.32 79.366.45 68.684.21 77.673.87 61.706.14 69.285.91 67.404.75 74.086.33
BBoxes 76.204.89 81.465.74 56.003.67 65.486.58 31.603.92 37.754.01 64.825.43 69.806.27
Points + BBoxes 75.965.11 80.593.86 69.034.32 78.066.02 38.975.38 46.676.49 72.705.07 79.754.73
MedSAM ViT-B BBoxes 65.189.45 69.229.51 68.917.54 75.127.60 74.304.94 77.814.82 63.697.40 66.107.28
Table 6: Comparison between different SAM pre-trained models and visual prompting techniques (meanstd)
4.3.4 SAM and Visual Prompting Comparison
Table 6 compares different SAM models and visual prompting techniques. We see that bounding boxes generally provided the best segmentation performance, except in Lung X-rays, where adding point prompts enhanced results. On the other hand, point prompts alone performed worse while combining points and bounding boxes slightly improved the results in some tasks, such as Lung X-ray (76.22 DSC, 81.30 NSD). In addition, the comparison of SAM, MedSAM, and SAM-Med2D demonstrates that SAM, despite not being pre-trained on medical data, performs well with bounding box prompts, achieving high scores in most modalities/tasks, including Lung CT. SAM-Med2D excels in fine-scaled segmentation, but struggles with larger structures, like lung lobes, where MedSAM performs better. The superior performance of SAM may be attributed to its use of a larger model architecture (ViT-H) compared to MedSAM and SAM-Med2D, which only offer ViT-B configurations.
@@ 4.4 Qualitative Segmentation Results
Lastly, we present qualitative segmentation results across the four imaging modalities evaluated in Fig. 2. Our proposed MedCLIP-SAMv2 consistently produced high-quality segmentation masks in weakly supervised settings. For all datasets except Brain MRI, the initial coarse segmentation was suboptimal. However, it provided a sufficient starting point for the zero-shot approach to refine coarse activation maps. For breast and brain tumors, the zero-shot results were notably better than those for Lung CT and Lung X-ray. In Lung CT, the primary challenge for the algorithm was distinguishing between the two lobes. The post-processed results showed one large, connected contour in the center. The zero-shot refinement slightly separated these two regions, though some artifacts persisted. However, the weakly supervised training effectively corrected these false activations, producing a high-quality segmentation map. For Lung X-ray, while the weakly supervised training improved upon the less precise zero-shot masks, the improvement was not as substantial as with Lung CT. Furthermore, we also included uncertainty maps for all predictions. For Brain MRI, high uncertainty was observed only at the edges of the segmentation, which is typical. For Breast Ultrasound, high uncertainty was observed at the borders of the segmentation, while the surrounding area outside the borders showed low uncertainty. In contrast, for Lung X-rays, slight uncertainty appeared in the center of the mask, increasing towards the edges. In the case of Lung CT, high uncertainty was observed both at the edges and in the center of the lung lobes. This was largely due to the artifacts present in the zero-shot pseudo-labels.
@@ 5 Discussion
The presented MedCLIP-SAMv2 framework demonstrates superior performance in zero-shot and weakly supervised medical image segmentation tasks across four critical medical imaging modalities: CT, MRI, Ultrasound, and X-ray. By leveraging BiomedCLIP and SAM with visual and text prompts, our method exhibits robust generalization, excelling in complex tasks such as brain and breast tumor segmentation, where smaller and intricate anatomical details pose challenges. This approach notably surpasses other SOTA zero-shot and few-shot methods, especially in difficult segmentation scenarios (see Table 1). Recent methods like [56] have demonstrated the potential of CLIP for zero-shot segmentation by decoupling the pixel-level and image-level classification tasks. However, such methods require fully supervised segmentation ground truth, limiting their application in settings where labels are scarce or noisy. In contrast, MedCLIP-SAMv2 bypasses this requirement and operates without relying on segmentation labels during training, offering a more scalable approach for medical imaging, particularly in weakly supervised settings.
@@
One of the key strengths of our framework lies in the integration of the novel M2IB, which effectively extracts meaningful information from medical images and texts, enhancing segmentation performance. The introduction of the DHN-NCE loss played a crucial role in fine-tuning BiomedCLIP, enabling the model to focus on challenging details while maintaining high performance across all tasks and modalities. Importantly, the combination of M2IB and DHN-NCE allowed the model to generate coarse segmentation masks that are later refined via SAM in a zero-shot setting, proving the versatility of the method without the need for ground truth annotations. The effectiveness of prompt design was another critical insight. Contextually rich, descriptive prompts yielded better results in complex tasks like tumor segmentation, where finer anatomical understanding is required. Conversely, more generic prompts sufficed for simpler tasks like lung segmentation, where larger, distinct structures allowed the model to achieve strong performance with less specific guidance.
@@
Our framework’s ability to operate in a weakly supervised paradigm further strengthens its potential clinical applicability. By using pseudo-labels from zero-shot segmentation to fine-tune the model, we saw notable improvements, particularly in lung CT segmentation, where the combination of zero-shot labels and weak supervision generated significant accuracy gains. To the best of our knowledge, we are the first to integrate uncertainty estimation through nnUNet with checkpoint ensembling by training on pseudo-segmentation data, providing a robust method for enhancing segmentation quality while offering insights into prediction confidence. Uncertainty measures are essential in clinical adoption, as they help identify regions where the model’s predictions are less certain, enabling clinicians to focus on areas that may require further examination or validation.
@@
Despite the original SAM model not being pre-trained on medical images, it showed strong performance in zero-shot settings, outperforming MedSAM and SAM-Med2D when provided with imperfect visual prompts like points or bounding boxes. This underscores the robustness of SAM to sub-optimal input conditions. Specifically, this can be seen in Fig. 2 where even suboptimal coarse segmentations can be refined using both zero-shot and weakly supervised methods. Looking ahead, future work will focus on extending our framework to handle 3D medical data, a crucial step in advancing the segmentation of volumetric imaging modalities like MRI and CT. Incorporating 3D models will enable our framework to better capture complex anatomical structures, further enhancing its clinical utility. Overall, our findings show that MedCLIP-SAMv2, with its novel integrated components, marks a significant step forward in the development of universal, interactive medical image segmentation. The framework’s adaptability across different tasks and its ability to operate with minimal labeled data emphasize its potential for clinical adoption, particularly in resource-constrained environments.
6 Conclusions
@@
We presented MedCLIP-SAMv2, a refined version of the original MedCLIP-SAM framework, improving segmentation performance with minimal supervision across CT, X-ray, Ultrasound, and MRI. By introducing the novel DHN-NCE loss for fine-tuning BiomedCLIP and leveraging SAM, our model achieved enhanced accuracy, particularly in complex tasks. MedCLIP-SAMv2 outperforms its predecessor through superior generalization and refined segmentation, demonstrating strong potential for clinical use in data-limited environments.

### v4

@@ 4.3.1 Effect of text prompt designs
We conducted a series of experiments to analyze the impact of various text prompt designs on zero-shot segmentation performance. In particular, we compared six different prompt configurations: P0 and P1 include the class name of the object to be segmented, while P2 and P3 consist of longer, descriptive single prompts, and finally P4 and P5 are ensembles of 20 text prompts. Note that P0, P2, and P4 are generic text prompts, while P1, P3, and P5 are more nuanced with subtypes of the target object of interests. For example, for Breast Ultrasound, P0 is “breast tumor” while P1 can either be “malignant breast tumor” or “benign breast tumor” depending on the tumor class. For P2, we used one descriptive sentence, such as “A medical breast mammogram showing a suspicious, irregularly shaped mass suggestive of a breast tumor.P3, on the other hand, includes descriptive text about a specific tumor subtype, like “A medical breast mammogram showing an irregularly shaped, spiculated mass suggestive of a malignant breast tumor.P4 and P5 are similar to P2 and P3, but they use an ensemble approach by averaging the text embeddings of 20 different prompts. Here, all descriptive clinical prompts are generated using GPT-4 (Achiam et al., 2023). For Lung CT, we evaluated solely on generic prompts as there is only one class available. As shown in Table 3, the choice of text prompt significantly influences segmentation performance. Class-specific prompts (P3) generally yielded better results for smaller structures like breast and brain tumors whereas generic prompts (P0, P2) performed better for larger structures like lungs in X-ray and CT scans, where simpler, more generic descriptions allowed the model to focus on larger areas. The best prompt configuration for each task is used to generate the results presented in Table 1.
Figure 4: Qualitative comparison of segmentation results. Coarse=post-processed saliency map, WSS=Weakly Supervised Segmentation and GT=Ground Truth. The uncertainty map corresponds to the weakly supervised segmentation.
4.3.2 Ablation Analysis of Algorithm Components
@@
Table 4 shows the contribution of each component of our framework in improving the average segmentation performance on all datasets. Starting with saliency maps generated using the M2IB, we achieved a baseline DSC of 46.23% and an NSD of 50.50%, providing an initial focus on key regions of interest. Fine-tuning BiomedCLIP with the proposed DHN-NCE loss raised the DSC to 49.10% and the NSD to 53.54%. Post-processing the saliency maps further enhanced the segmentation quality, allowing the model to better delineate foreground and background areas by refining the initial segmentation boundaries. Incorporating a connected component analysis step greatly impacted the results, increasing the DSC to 57.89% and the NSD to 61.54%, as it eliminated small, irrelevant clusters and reduced noise, improving overall precision. With the integration of SAM and the use of visual prompts, such as bounding boxes or points, our zero-shot method yielded a substantial improvement, achieving a DSC of 77.61% and an NSD of 81.56%. Finally, weakly supervised training with checkpoint ensembling further refined the segmentation quality by leveraging pseudo-labels generated from the zero-shot method. By using these pseudo-labels to fine-tune a segmentation network, we were able to reach a final DSC of 82.11% and an NSD of 87.33%.
@@
1: Saliency Maps 46.238.58 50.508.86
2: + DHN-NCE Fine-tuning 49.108.46 53.548.62
3: + Post-processing 51.627.57 55.237.47
4: + Connected Component Analysis 57.897.87 61.548.02
5: + SAM 77.616.82 81.567.00
6: + nnUNet Ensemble 82.118.49 87.338.46
Table 4: Effect of different components (%, meanstd)
4.3.3 Impact of Saliency Maps Generation Methods
As shown in Table 5, M2IB achieved the highest performance across all tasks, with an average DSC of 77.61% and NSD of 81.56% when using the fine-tuned BiomedCLIP model. In both its pre-trained and fine-tuned forms, M2IB significantly outperformed gScoreCAM and GradCAM ($p$ $<$ 0.05). BiomedCLIP fine-tuning improved the scores across all saliency map techniques on average, with the largest gains seen in M2IB, which improved by 3.92% in DSC and 4.24% in NSD compared to its pre-trained version.
@@ 4.3.3 Impact of Saliency Maps Generation Methods
Pre-trained BiomedCLIP M2IB 73.697.58 77.327.43
gScoreCAM 58.926.67 62.196.02
GradCAM 29.218.74 31.368.44
Fine-tuned BiomedCLIP M2IB 77.616.82 81.567.00
gScoreCAM 60.526.41 63.896.39
GradCAM 30.118.92 32.618.83
Table 5: Comparison between different Saliency Map techniques as well as the pre-trained and fine-tuned BiomedCLIP on the overall performance (%, meanstd)
@@
SAM ViT-H Points 65.569.89 68.209.97 65.548.45 70.738.32 75.793.44 80.883.52 61.496.25 63.906.74
BBoxes 77.769.52 81.119.89 76.527.06 82.237.12 70.555.38 74.125.77 80.385.81 82.035.94
Points + BBoxes 74.3810.57 79.6010.62 75.488.66 80.298.64 73.305.94 79.226.12 62.836.72 64.576.99
SAM-Med2D ViT-B Points 73.129.51 75.169.13 66.789.97 70.129.75 60.587.43 64.427.73 65.947.17 68.057.99
BBoxes 75.2210.04 80.0310.94 55.219.85 61.349.93 30.1811.15 36.3511.23 63.108.57 68.598.48
Points + BBoxes 74.8310.78 79.5010.12 67.8510.96 72.0410.45 37.238.69 44.909.37 71.228.09 78.058.11
MedSAM ViT-B BBoxes 63.5011.42 68.1111.25 67.6812.75 73.8912.67 73.036.03 76.236.02 62.147.80 65.007.11
Table 6: Comparison between different SAM pre-trained models and visual prompting techniques (%, meanstd)
4.3.4 Comparison of Visual Prompts for SAM
Table 6 compares different SAM models and visual prompting techniques. We see that bounding boxes generally provided the best segmentation performance, except in Lung X-rays, where adding point prompts enhanced results. On the other hand, point prompts alone performed worse except in certain tasks, such as Lung X-ray (75.79% DSC, 80.88% NSD). In addition, the comparison of SAM, MedSAM, and SAM-Med2D demonstrates that SAM, despite not being pre-trained on medical data, performs well with bounding box prompts, achieving high scores in most modalities/tasks, including Lung CT. SAM-Med2D excels in fine-scaled segmentation, but struggles with larger structures, like lung lobes, where MedSAM performs better. The superior performance of SAM may be attributed to its use of a larger model architecture (ViT-H) compared to MedSAM and SAM-Med2D, which only offer ViT-B configurations.
@@ 4.4 Qualitative Segmentation Results
Lastly, we present qualitative segmentation results across the four imaging modalities evaluated for our proposed method in Fig. 4. Our proposed MedCLIP-SAMv2 consistently produced high-quality segmentation masks in weakly supervised settings. For all datasets except Brain MRI, the initial coarse segmentation was suboptimal. However, it provided a sufficient starting point for the zero-shot approach to refine coarse activation maps. For breast and brain tumors, the zero-shot results were notably better than those for Lung CT and Lung X-ray. In Lung CT, the primary challenge for the algorithm was distinguishing between the two lobes. The post-processed results showed one large, connected contour in the center. The zero-shot refinement slightly separated these two regions, though some artifacts persisted. However, the weakly supervised training effectively corrected these false activations, producing a high-quality segmentation map. For Lung X-ray, while the weakly supervised training improved upon the less precise zero-shot masks, the improvement was not as substantial as with Lung CT. Furthermore, we also included uncertainty maps for all predictions. For Brain MRI, high uncertainty was observed only at the edges of the segmentation, which is typical. For Breast Ultrasound, high uncertainty was observed at the borders of the segmentation, while the surrounding area outside the borders showed low uncertainty. In contrast, for Lung X-rays, slight uncertainty appeared in the center of the mask, increasing towards the edges. In the case of Lung CT, high uncertainty was observed both at the edges and in the center of the lung lobes. This was largely due to the artifacts present in the zero-shot pseudo-labels.
@@ 5 Discussion
The proposed MedCLIP-SAMv2 framework demonstrates superior performance in zero-shot and weakly supervised medical image segmentation tasks than SOTA methods and the original MedCLIP-SAM method (Koleilat et al., 2024b) across four critical medical imaging modalities (CT, MRI, Ultrasound, and X-ray). By leveraging BiomedCLIP and SAM with text and visual prompts, our method exhibits robust domain and task generalization, excelling in complex tasks, such as brain and breast tumor segmentation, where smaller and intricate anatomical details pose challenges in typical segmentation tasks. Our approach notably surpasses other SOTA zero-shot and few-shot methods, especially in difficult segmentation scenarios (see Table 1). Recent methods like (Ding et al., 2022) have demonstrated the potential of CLIP for zero-shot segmentation by decoupling the pixel-level and image-level classification tasks in natural vision applications. However, such methods require fully supervised segmentation ground truths, limiting their application in settings where labels are scarce or noisy, like medical image segmentation. In contrast, MedCLIP-SAMv2 bypasses this requirement and operates without relying on segmentation labels during training, offering a more scalable approach for medical imaging, particularly in weakly supervised settings.
@@
Compared with the original MedSAM-CLIP, the component updates in MedSAM-CLIPv2 have greatly contributed to the performance improvement. One of the key strengths of our framework lies in the integration of M2IB for radiological tasks, which effectively extracts meaningful information from medical images and texts, enhancing segmentation performance. The introduction of the DHN-NCE loss played a crucial role in fine-tuning BiomedCLIP, enabling the model to focus on challenging details while maintaining high performance across all tasks and modalities. Importantly, the combination of M2IB and DHN-NCE allowed the model to generate coarse segmentation masks that are later refined via SAM in a zero-shot setting (see Table 5), proving the versatility of the method without the need for ground truth annotations. Finally, the effectiveness of prompt design was another critical insight. Contextually rich, descriptive prompts yielded better results in complex tasks like tumor segmentation, where finer anatomical understanding is required. Conversely, more generic prompts sufficed for simpler tasks like lung segmentation, where larger, distinct structures allowed the model to achieve strong performance with less specific guidance. This insight suggests the importance of tailoring the text prompts in visional language models for specific radiological tasks. This contrasts with findings from other studies that used the frozen BiomedCLIP encoder with an added decoder head for segmentation transfer learning, where text prompts had little impact on segmentation quality (Poudel et al., 2023). The choice of BiomedCLIP over CLIP also facilitates the success of our method. Figure 5 shows the latent representations produced by CLIP and BiomedCLIP (both utilizing the same architecture i.e. ViT-B/16) of sample medical images. The latter shows that the BiomedCLIP model learns to encode meaningful latent representations of salient regions within medical scans from only natural language supervision, facilitating its ability to highlight disease-relevant regions across various modalities compared to CLIP where the subtle visual cues found in medical images are not sufficiently captured or distinguished.
Figure 5: Diagram showing upsampled feature representations from the last transformer layer of CLIP and BiomedCLIP. Feature Maps were upsampled using FeatUp (Fu et al., 2024) for visualization purposes.
@@
Our framework’s ability to operate in a weakly supervised paradigm further strengthens its potential clinical applicability. By using pseudo-labels from zero-shot segmentation to fine-tune the model, we observed notable improvements, particularly in lung CT segmentation, where the combination of zero-shot labels and weak supervision generated significant accuracy gains. To the best of our knowledge, we are the first to integrate uncertainty estimation through nnUNet with checkpoint ensembling by training on pseudo-segmentation data, providing a robust method for enhancing segmentation quality while offering insights into prediction confidence for potential end users. Uncertainty measures are essential in clinical adoption, as they help identify regions, where the model’s predictions are less certain, enabling clinicians to focus on areas that may require further examination or validation.
@@
Despite the original SAM model not being pre-trained on medical images, it showed strong performance in zero-shot settings, outperforming MedSAM and SAM-Med2D when provided with imperfect visual prompts like points and/or bounding boxes. This underscores the robustness of SAM to suboptimal input conditions as highlighted by (Huang et al., 2024). Specifically, this can be seen in Fig. 4, where even coarse segmentations can be refined using both zero-shot and weakly supervised methods. Looking ahead, our future work will focus on extending our framework to handle 3D medical data, a crucial step in advancing the segmentation of volumetric imaging modalities like MRI and CT. Incorporating 3D models will enable our framework to better capture complex anatomical structures, further enhancing its clinical utility. Overall, our findings show that MedCLIP-SAMv2, with its integrated components, marks a significant step forward in the development of universal, interactive medical image segmentation. The framework’s adaptability across different tasks and its ability to operate with minimal labeled data emphasize its potential for clinical adoption, particularly in resource-constrained settings. For our exploration, we focused on radiological tasks, with image modalities having more distinct characteristics than natural images. In the future, we will further incorporate and assess the performance of photograph-based biomedical images, such as histopathological images and surgical video with our proposed framework.
6 Conclusion
@@
We presented MedCLIP-SAMv2, an upgraded version of the original MedCLIP-SAM framework, significantly improving segmentation performance with minimal supervision across CT, X-ray, Ultrasound, and MRI. By introducing the novel DHN-NCE loss for fine-tuning BiomedCLIP and leveraging SAM, our model achieved enhanced accuracy, particularly in complex tasks. MedCLIP-SAMv2 outperforms its predecessor through superior generalization and refined segmentation, demonstrating strong potential for clinical use in data-limited environments.